OpenStudy (anonymous):

how do you simplify: 2 radical 50- 3 radical 8 I don't understand how to deal with the number inside of the radical symbol

OpenStudy (amistre64):

undo it

OpenStudy (amistre64):

8 factors into a perfect square; so does 50

OpenStudy (anonymous):

factor the radical and look for squares

OpenStudy (anonymous):

how do you factor the radical tho. They confuze me

OpenStudy (amistre64):

1 4 9 16 25 36 49 look and see if you can factor out a perfect square

OpenStudy (anonymous):

2 3 4 5 6 and 7 i understand that

OpenStudy (amistre64):

4*2 = 8 --> 2sqrt(2) 25*2 = 50 --> 5sqrt(2)

OpenStudy (amistre64):

\[\sqrt{8}=\sqrt{4*2}=\sqrt{4}*\sqrt{2}=2\sqrt{2}\]

OpenStudy (anonymous):